OpenStudy (anonymous):

I need help with # 3 math prob will some body help plz

OpenStudy (anonymous):

2. find the height of a tringel with area= 110 ft and base =10ft 3.if the area of parallelogram ABCD = 246 mm and h= 20.5 mm what is the length of the base of tringle ABD

OpenStudy (texaschic101):

Area of a triangle can be found with this formula : A = 1/2bh so what is the height of the first triangle ? what is the base ?

OpenStudy (anonymous):

the parallelogram is for #3

OpenStudy (anonymous):

oo ya I did and I got 21 I wanted to make sure it was rigth

OpenStudy (texaschic101):

do you know the height of your first triangle ?

OpenStudy (anonymous):

7in

OpenStudy (texaschic101):

A = (7 x 6) / 2 = 42/2 = 21 you are correct. Area of triangle is 21

OpenStudy (anonymous):

ya ok and what about the other ones

OpenStudy (texaschic101):

2. A = 1/2bh A = 110 b = 10 now we sub 110 = 1/2(10)h 110 = 5h 110/5 = h 22 = h check.. A = 1/2bh 110 = 1/2(10)(22) 110 = 5(22) 110 = 110 (correct) height is 22 ft

OpenStudy (anonymous):

yup got it right :)

OpenStudy (texaschic101):

I am sorry but I do not have time to do your last one...I have to go somewhere...sorry :(

OpenStudy (anonymous):

k thanks any way